# HANS-Net: Hyperbolic Convolution and Adaptive Temporal Attention for Accurate and Generalizable Liver and Tumor Segmentation in CT Imaging

### v1

@@ 7   Conclusion
We have developed and thoroughly evaluated HANS-Net, an innovative segmentation framework that achieved a remarkable Dice score in the segmentation of both the liver and tumors. The framework obtained Dice scores of 96.67% for the liver and 89.84% for tumors, and IoU values of 93.79% for the liver and 82.39% for tumors, respectively. The model was advanced in architecture, containing six modules designed to boost and enhance segmentation performance. The dataset was further assessed on the 3D-IRCADb-01 dataset to demonstrate its robust generalization capabilities. However, it faced a slight decline in tumor segmentation performance compared to the results obtained from the LiTS dataset.
The significance of this work lies in its integrated modules, which serve several key functions: accurately defining organ boundaries, extracting features from multiple levels, and quantifying prediction confidence. All of these contribute to the strong performance of the model. HANS-Net outperformed all state-of-the-art and pre-trained models in both liver and tumor segmentation tasks. These results confirm that HANS-Net is capable of achieving high segmentation performance, making it a promising tool for use in clinical and medical imaging applications.
While performance is promising, further work is needed to improve generalization to diverse imaging settings, including other modalities and multiorgan segmentation, as well as to enhance runtime efficiency for clinical deployment. By providing anatomically precise, uncertainty-aware segmentation, HANS-Net can aid radiologists in early detection and treatment planning and has the potential to improve clinical workflows.
@@ Acknowledgments
In this study, we used five publicly available datasets, including Liver Tumor Segmentation Challenge (LiTS) dataset [26], the 3D-IRCADb-01 dataset [27].

### v2

@@ 7 Conclusion
We develop and thoroughly evaluate HANS-Net, an innovative segmentation framework that achieves remarkable Dice scores for both liver and tumor segmentation. The key experimental findings and theoretical conclusions are summarized as follows:
  • HANS-Net attains high segmentation accuracy on the LiTS dataset, with Dice scores of 96.67% for the liver and 89.84% for tumors, and IoU values of 93.79% and 82.39%, respectively.
  • Cross-dataset evaluation on the AMOS 2022 dataset achieves Dice scores of 85.09% and IoU of 76.66% across 15 abdominal organs, demonstrating strong generalization ability across diverse anatomical structures and imaging variations.
  • The significance of this work lies in its six integrated modules, which collectively define organ boundaries, extract multi-level features, and quantify prediction confidence, supporting the robust performance of the model.
Future work will focus on enhancing generalization across diverse imaging modalities while optimizing runtime efficiency for clinical deployment. While our current evaluation focuses on standard slice-level segmentation metrics, including Dice, IoU, ASSD, and VOE, future work should consider reporting these metrics at the volumetric level to maintain consistency with the original LiTS benchmark protocol. Additionally, while the model operates on resized 2D slices, it employs dynamic sequential weighting to preserve volumetric continuity, a mechanism that can be further adapted to explicitly enable 3D reconstruction capabilities.
@@ Acknowledgments
In this study, we used two publicly available datasets, including Liver Tumor Segmentation Challenge (LiTS) dataset [35], and AMOS 2022 dataset[36].